A man has a lot with a frontage of 26 yds 2 ft. He wishes to set 5 shade trees along the front, equally spaced, one at each end with three between. How far apart will they be?
Answer by stanbon(75887)

A man has a lot with a frontage of 26 yds 2 ft. He wishes to set 5 shade trees along the front, equally spaced, one at each end with three between. How far apart will they be?
---------
26rds 2 ft = 26*3+2 ft = 80 ft
----
Placing 3 evenly between the ends makes 4 even intervals.
Ans:: 80/4 = 20 ft = 6 yds 2 ft
=================
Cheers,
Stan H.
----------------